all the black face cards are removed from a pack of 52 playing cards. the remaining cards are well shuffld and then a card is drawn at random. find the probability of getting a
1. face card
2. red card
3. black card
4. king

Found 2 solutions by ewatrrr, Edwin McCravy:
Answer by ewatrrr(24785)

 
Hi,
46 cards with 6 face cards remaining (all black face cards removed )
P(1 FC) = 6/46
P(red card) = 26/48
P(black card) = 20/48
P(K) = 2/48
reduce fractions

Answer by Edwin McCravy(20054)


all the black face cards are removed from a pack of 52 playing cards

A♥   2♥   3♥   4♥   5♥   6♥   7♥   8♥  9♥  10♥  J♥  Q♥  K♥ 
A♦   2♦   3♦   4♦   5♦   6♦   7♦   8♦  9♦  10♦  J♦  Q♦  K♦

 1. P(face card) = 6 out of 26 = 6/26 = 3/13 
 2. P(red card) = 26 out of 26 = 26/26 = 1
 3. P(black card) = 0 out of 26 = 0/26 = 0
 4. P(king) = 2 out of 26 = 2/26 = 1/13

Edwin